From PREX Wiki
Jump to: navigation, search

Back to Main Page >> Analyzer_Meeting

previous meeting << >> following meeting

Logistic information

Meeting room at JLab: CC L210A (entrance is in the lobby across from the 2nd floor auditorium entrance)

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  419 045 306
 Bluejeans link: https://bluejeans.com/419045306


  1. respin of APEX runs: runlist
    • What output do we have?
      • TY: high current with SAMs on have been respun for 50 and 40 uA runs.
      • TY: right now we look at the noise for BPMs and SAMs.
      • KP: would you know if all the events are processed? TY: will figure out how to best do that. SR: if the prompt summary exists then it finished.
      • KP: the prompt summary text output should also be investigated.
      • TY: we used the ds3x bcm to normalize the SAMs
      • CG: we should take a look at Caryn's panguin file. TY: will do
    • Agregator?
      • CG: SR and CG will take a look at Cameron's code.
    • Pair tree plots?
      • TY: working on pair tree plots and post-pan regression with pairs.
    • Do we have a "process a run" analysis chain ready? Do we want to set it up as a script? >> This will develop into the prompt analysis script.
      • TY: we have a basic script. PK: there are Qweak examples for watching files in the directories that we copied over to adaq.
      • KP: the shift crew should check the basic quality of the data output.
  2. "Operations" branch
    • PK: Tao updated files and committed and I pushed them to the repo.
  3. regression analysis
    • PK: we are pretty close to having a working copy of the data handler. Still on track to have a regression analysis working by the end of April.
  4. feedback
    • KP: could feedback hang off the data handler? PK: Amali should have a discussion with Paul, because there should not be that much restructuring to implement feedback.
  5. Beam modulation
    • CG: will follow up with Victoria.
  6. VQWK adcs
    • CG: all working.
  7. EPICS run-start and fast logger (and EPICS decoder to slow tree)
    • RR: updated both of them (removed the non-working EPICS variables). Organized the fastLogger.
  8. bmwClient has issues with not pausing when CODA isn't in progress, being stopped and restarted, and to have parameters changed. This affects feedback as well and can change values when a CODA run is started with the gateway open.
    • PK: the bmwClient doesn't get data from CODA. My intention is to add variables to the BMW client that are written at pre-start, end of run, and download.
  9. Bob's DAQ tasklist
    • Timing: mostly understood, still small mysteries (LNE)
      • PK: we restructured the timing between the injector and CH crates.
    • Synchronization-check system. w/help from Ed Jastrzembski
      • Cameron working on.
    • Clean up cabling, labels, fasteners - First need to examine all signals
      • CH crate has a preliminary cleanup and Caryn is doing
    • Clean up CRL -- eliminate so-called "dirty trick"
      •  ??
    • Check the 7200 scaler setup at injector, find a spare 7200.
      • Talked to Bob related to timing. Event counters
    • Pedestal studies, aka cross-talk
      • Tao will look into it soon.
    • Green Monster rework: Adam.
      • AZ: having a separate issue from bmw.
    • automatic production of run-dependent database
      • Sanghwa is working on this.
    • alarms like end-of-run voice
      • later
    • e-logbook entries
      • later
    • Document everything. Circuit diagrams, cable map, "how to".


Tao Y, Ciprian G, Adam Z, Kent P, Sakib R, Robert R, Devi A